# Feature Alignment: Rethinking Efficient Active Learning via Proxy in the Context of Pre-trained Models

### v1

@@ Abstract
Fine-tuning the pre-trained model with active learning holds promise for reducing annotation costs. However, this combination introduces significant computational costs, particularly with the growing scale of pre-trained models. Recent research has proposed proxy-based active learning, which pre-computes features to reduce computational costs. Yet, this approach often incurs a significant loss in active learning performance, which may even outweigh the computational cost savings. In this paper, we argue the performance drop stems not only from pre-computed features’ inability to distinguish between categories of labeled samples, resulting in the selection of redundant samples but also from the tendency to compromise valuable pre-trained information when fine-tuning with samples selected through the proxy model. To address this issue, we propose a novel method called aligned selection via proxy to update pre-computed features while selecting a proper training method to inherit valuable pre-training information. Extensive experiments validate that our method significantly improves the total cost of efficient active learning while maintaining computational efficiency.
@@ 1 Introduction
Training effective deep neural networks typically requires large-scale data [11, 16, 23]. However, the high cost of acquiring data, especially annotated data, poses a significant challenge for practitioners [3, 25]. Active learning and the pretraining-finetuning paradigm are widely adopted strategies to address this challenge. Active learning reduces the demand for extensive annotation by iteratively selecting and labeling the most informative samples [28]. Additionally, the pretraining-finetuning method leverages large-scale unsupervised pre-training to create a powerful foundational model [7, 14, 5, 17], enabling exceptional performance on downstream tasks with only a limited amount of labeled data. To further minimize labels’ demands, researchers have naturally turned their attention to active fine-tuning, which fine-tunes the pre-training model with labeled samples selected by active learning strategies [37, 6, 2]. However, as the scale of pre-training models continues to grow [8], the training cost of active fine-tuning, which often is overlooked in traditional active learning, becomes a challenge [10].
@@
Figure 1: When achieving the accuracy equivalent to randomly selecting 100k samples through margin active learning (selecting 10k for each active learning iteration), we compared the labeling cost savings and the increased active learning training cost between standard active learning pipelines and the efficient active learning method, SVPp. We used a resnet50 model, and the training cost was priced based on AWS EC2 P3 instances (following existing paper [42]), while annotation costs were estimated using AWS Mechanical Turk [^1] https://aws.amazon.com/sagemaker/groundtruth/pricing/ with triple reviews.
@@
To address this issue, recent research has introduced an efficient active learning framework known as Selection via Proxy based on pre-trained features (SVPp) [42]. SVPp begins by forwarding the entire dataset through a pre-trained model once, recording pre-trained features for all samples. Subsequently, in multiple iterations of active learning, a simple MLP classifier is trained using the pre-computed features for sample selection. After the sample selection, the entire model is fine-tuned once. However, we observe that, in comparison to the standard active learning method (fine-tuning the whole pre-trained model in each active learning iteration), SVPp may compromise the effectiveness of active learning, resulting in additional annotation costs. As depicted in fig. 1, we noticed that while SVPp significantly reduces active learning time (training costs), its decline in active learning performance leads to a notable increase in annotation costs. Practitioners therefore have to face the hassle of weighing computational efficiency against overall cost.
@@
In light of this challenge, in this paper, we analyze the factors contributing to the decline in active learning performance when using SVPp in sec. 4. We initially observe an intuition-aligned phenomenon: with an increase in the number of labels, fine-tuned features become more capable of distinguishing sample categories than pre-trained features. Consequently, this leads to SVPp selecting redundant samples, where redundancy refers to SVPp selecting the samples that the fine-tuned model can already predict correctly. In addition, we make an intriguing discovery: during the fine-tuning, samples selected by SVPp tend to induce more significant modifications to the pre-trained model, resulting in a decreased performance of the fine-tuned model.
@@
Based on these insights, we propose an aligned selection via proxy strategy, ASVP, to enhance the performance of efficient active learning in sec. 5. First, we align the pre-computed features utilized by the proxy model with the one used in standard active learning. Specifically, we update the pre-computed features based on the fine-tuned model when the pre-trained features’ ability to distinguish sample categories significantly lags behind the fine-tuned features. In addition, we switch the training method between linear-probing then fine-tuning (LP-FT) [20] and fine-tuning (FT) to mitigate the impact of samples selected by the proxy model on the model’s performance.
@@
Extensive experiments validate that our method notably improves the effectiveness of efficient active learning, achieving comparable or better total cost savings comparable to those of standard active learning, while still maintaining computational efficiency.
@@
The contributions of this paper can be summarized as follows: (1) We empirically analyze the factors contributing to the performance drop when employing SVPp. The proxy model picks redundant samples where fine-tuned features outperform pre-trained ones. Also, it overlooks samples that are crucial to retain valuable pre-trained model information. (2) We propose a novel efficient active learning approach, aligned selection via proxy (ASVP), based on the above analysis. It improves the performance of efficient active learning while incurring minimal computation time. (3) We introduce a novel and practical evaluation metric for efficient active learning: the sample saving ratio. This metric directly quantifies the savings in annotation achieved by employing active learning strategies compared with the random baseline. It also facilitates the estimation of overall savings, including savings in both computational and labeling costs, thereby assisting practitioners in making a decision on whether to use efficient active learning strategies.
@@ 2 Related Work
Active learning is a technique that minimizes the number of labels needed for model training by selectively requesting annotations for samples. Numerous active learning strategies have been proposed, primarily encompassing uncertainty sampling [21, 30, 13, 18, 36], diversity sampling (feature space coverage) [31, 24], their combinations [38, 1, 32], and some learning-based approaches [33, 40, 34]. With the advancement of large-scale unsupervised pre-training and fine-tuning, an increasing number of researchers have turned their attention to active fine-tuning, which fine-tuning the pre-trained model with samples selected by active learning strategies. Given the success of the pre-training fine-tuning mode, most works of active fine-tuning have focused on developing active learning strategies that can work with extremely limited annotations [15, 39, 35]. Furthermore, researchers have validated the performance of existing active learning methods when applied to pre-training fine-tuning mode, demonstrating that traditional margin sampling exhibits strong performance [42, 12].
Efficient active learning. Most active learning methods operate in a multi-round mode, iteratively training models and selecting samples to label. This mode entails significant training time and costs. In the context of active fine-tuning, this phenomenon becomes more pronounced, as larger models are better equipped to effectively leverage data in unsupervised pre-training [8, 26].
Increasing the active learning batch size, which refers to the number of samples selected per active learning iteration [9, 43], is one solution to enhance computational efficiency. However, it weakens the performance of active learning. Moreover, even when increasing the active learning batch size, there remains the considerable computational cost associated with training the entire network multiple times. Therefore, recent research has explored single-shot active learning based on pre-trained features [37]. However, as the number of labels increases, the effectiveness of this method gradually diminishes in comparison to other standard active learning strategies.
Another solution to boost active learning efficiency is to use less computationally intensive models or training methods as proxy tasks for sample selection, known as Selection via Proxy (SVP) [10]. Typical proxy tasks include reducing model complexity, such as using models like ResNet8 or ResNet14, early stopping, or training linear classifiers or MLPs based on pre-trained features [42]. However, these proxy tasks often have an impact on active learning performance and may result in savings in computation costs not outweighing the increase in annotation costs. As a result, practitioners are often faced with a trade-off between computational efficiency and the overall cost.
@@ 3 Preliminary
We briefly review the Selection via Proxy framework based on pre-computed features, SVPp, as shown in fig. 2. We denote the labeled set as $L$, the unlabeled set as $U$. The pre-trained neural network backbone is defined as $f(\cdot;\theta_{p}):\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $\theta_{p}$ represents weights of the pre-trained model, $\mathcal{X}$ is the data space and $\mathbb{R}^{d}$ is normalized feature space. And we define the predictor as $h(\cdot;\theta_{h}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, where $\theta_{h}$ is weights of the predictor and $\mathbb{R}^{c}$ is the output space. SVPp follows a three-step process. First, all samples are forward-passed through $f(\cdot;\theta_{p})$, and the pre-trained features are saved. Subsequently, the proxy model is trained, and samples are selected iteratively based on the active learning strategy. In this process, the predictor $h(\cdot;\theta_{h})$ serves as the proxy model, utilizing pre-computed features as its input. The final step is fine-tuning the pre-trained model $h(f(\cdot;\theta_{p});\theta_{h})$ based on the obtained labels.
@@ 4 Observation and Analysis
To analyze the factors contributing to the decline in SVPp active learning performance, we investigated the differences in sample selection between the fine-tuned model, $h(f(\cdot;\theta_{p});\theta_{h})$, and the proxy model, $h(\cdot;\theta_{h})$. Our analysis is based on uncertainty sampling because uncertainty-based methods are widely employed in active learning strategies. Recent benchmarks have demonstrated uncertainty sampling as the state-of-the-art approach when fine-tuning pre-trained model [42, 12]. As illustrated in fig. 3, compared to the samples selected by the fine-tuned model, the proxy model misses samples from regions A and B while additionally selecting samples from regions C and D. Building on the rationale that adding samples already correctly predicted by the model seldom improves its performance, whereas incorporating samples the model predicts incorrectly often leads to performance enhancements. The impact of the proxy model on active learning performance primarily stems from missing samples in region B (missing region) and selecting samples redundantly in region C (redundant region).
@@
Figure 3: Sample Selection Discrepancy: Proxy vs. Fine-tuned Models. The two axes represent the confidence of predictions for the proxy model and the fine-tuned model, with the positive half-axis indicating correct predictions and the negative half-axis indicating incorrect predictions. Uncertainty-based active learning selects samples with the lowest prediction confidence, meaning the proxy model chooses samples from regions O, C, and D, while the fine-tuned model selects samples from regions O, A, and B. For convenience in further analysis, we denote the incorrect predictions made by the proxy model in region B as B1, and the correct predictions as B2. Both the proxy model and the fine-tuned model confidently predict the remaining white region, hence they do not select samples from it.
(a) Active Learning Accuracy
(b) Exchange-head
Figure 4: Comparision of exchange-head accuracy and active learning accuracy for samples selected via fine-tuned model (standard), proxy model, and replace experiment (replacing samples selected by the proxy model with those from the missing region).
Observation 1: The absence of samples from the missing region leads to a decline in the performance of the fine-tuned model. We employed one of the uncertainty-based active learning strategies, margin sampling, in three ways on the ImageNet dataset: (1) using the fine-tuned model, (2) using a proxy model, and (3) a replacement experiment where the top 20% of the samples with the largest margins in the selection of the proxy model were replaced with samples from the region B2. Additionally, considering that replacing samples with smaller margins in the fine-tuned model could potentially affect the final accuracy, we conducted another replacement experiment, replacing to least margin. In this setup, we directly replaced the top 20% samples with the largest margins in the selection of the proxy model with samples of the smallest margin on the fine-tuned model (region A and region B). The accuracy of the fine-tuned model is illustrated in fig. 3(a). Replacing samples with missing region samples shows very close active learning accuracy to that achieved based on the fine-tuned model. However, simply replacing them with samples that have the smallest margin in the fine-tuned model does not improve model accuracy.
As illustrated in fig. 3, samples in region B2 are well-distinguished in the pre-trained feature space (predicted accurately and confidently by the proxy model). However, the fine-tuned model struggles to predict these samples correctly, displaying high uncertainty in its predictions. Therefore, we speculate that replacing the samples selected by the proxy model with samples from this region improves active learning performance for the following reasons:
Potential reason: The samples selected from the missing region B2 potentially aid the fine-tuned model in retaining valuable pre-trained feature information, thereby enhancing active learning performance. We conducted exchange-head experiments to determine how much the fine-tuned model alters the feature space of the pre-trained model. Specifically, we placed the fine-tuned classifier on the pre-trained backbone to evaluate accuracy. Higher accuracy indicates less alteration of the pre-trained model by fine-tuning  [29]. As shown in fig. 3(b), samples chosen by the proxy model induce the most changes in the pre-trained feature space, while replacing the 20% samples selected by the proxy model with missing region samples significantly reduces alterations to the pre-trained model.
Observation 2: As the number of annotations increases, the proxy model selects more samples from redundant regions. The experiments on CIFAR-10 as shown in fig. 4(a). This observation aligns with intuition: as annotations increase, the features of the fine-tuned model become more adept at distinguishing between sample categories than those of the proxy model. Consequently, there is a gradual increase in the percentage of redundant samples chosen by the proxy model in each round. This implies that, as annotations increase, the influence of sample selection from the redundant region becomes more pronounced in active learning performance.
(a) Ratio of Redundant Samples
(b) Active Learning Accuracy
Figure 5: Observation of redundant region samples. (a) the proportion of redundant region samples among the newly added samples of each active learning round, where 10 rounds of margin sampling on CIFAR-10, selecting 200 samples per round. (b) Active learning accuracy comparison among selecting margin samples via fine-tuning, selecting samples from the redundant region, and random sample selection.
Observation 3: The samples selected from the redundant region contribute little to improving the fine-tuned model. We conducted experiments on the CIFAR-10 dataset with margin sampling. The initial pool was set to 1000 samples, and subsequently, 1000 samples were added through (1) fine-tuning the model, (2) samples from the redundant region, and (3) random selection. The results are depicted in fig. 4(b). The accuracy of adding samples from redundant regions is almost the same as that of adding samples randomly and is lower than the results obtained by selecting samples using the fine-tuned model.

### v2

@@ Abstract
Fine-tuning the pre-trained model with active learning holds promise for reducing annotation costs. However, this combination introduces significant computational costs, particularly with the growing scale of pre-trained models. Recent research has proposed proxy-based active learning, which pre-computes features to reduce computational costs. Yet, this approach often incurs a significant loss in active learning performance, sometimes outweighing the computational cost savings. This paper demonstrates that not all sample selection differences result in performance degradation. Furthermore, we show that suitable training methods can mitigate the decline of active learning performance caused by certain selection discrepancies. Building upon detailed analysis, we propose a novel method, aligned selection via proxy, which improves proxy-based active learning performance by updating pre-computed features and selecting a proper training method. Extensive experiments validate that our method improves the total cost of efficient active learning while maintaining computational efficiency. The code is available at https://github.com/ZiTingW/asvp.
@@ 1 Introduction
Training effective deep neural networks typically requires large-scale data (Deng et al., 2009; He et al., 2016; Liu et al., 2021). However, the high cost of acquiring data, especially annotated data, poses a significant challenge for practitioners (Budd et al., 2021; Norouzzadeh et al., 2021). Active learning and the pre-training fine-tuning paradigm are widely adopted strategies to address this challenge. Active learning reduces the demand for extensive annotation by iteratively selecting and labeling the most informative samples (Ren et al., 2021). Additionally, the pre-training fine-tuning method leverages large-scale unsupervised pre-training to create a powerful foundational model, enabling exceptional performance on downstream tasks with only a limited amount of labeled data (Chen et al., 2020a; Grill et al., 2020; Caron et al., 2021; He et al., 2022). To further minimize labels’ demands, researchers have turned their attention to active fine-tuning, which fine-tunes the pre-training model with labeled samples selected by active learning strategies (Xie et al., 2023; Chan et al., 2021; Bengar et al., 2021). However, as the scale of pre-training models continues to grow, the sample selection time of active learning, which often is overlooked in traditional active learning, becomes a challenge (Coleman et al., 2019). As illustrated in fig. 2, larger pre-training models lead to higher accuracy, but also significantly increase the time required for active learning sample selection.
Figure 1: Comparing sample selection time and accuracy in active learning (margin sampling) with different scale models. The active learning included a total of 16 iterations, selecting a total of 400k samples on the ImageNet dataset. All models are pre-trained using BYOL-EMAN (Cai et al., 2021).
@@
Figure 2: Comparison of the labeling and training costs across random sampling, the standard active learning pipeline, the efficient active learning method (SVPp), and our method (ASVP), using margin sampling to achieve the same accuracy as randomly selecting 400k samples on the ImageNet dataset. A ResNet-50 model was employed, with training costs calculated based on AWS EC2 P3 instances (following Zhang et al., 2024), and annotation costs estimated using AWS Mechanical Turk [^1] https://aws.amazon.com/sagemaker/groundtruth/pricing/ with triple reviews.
@@
To address this issue, recent research has introduced an efficient active learning framework known as Selection via Proxy based on pre-trained features (SVPp) (Zhang et al., 2024). SVPp begins by forwarding the entire dataset through a pre-trained model once, recording pre-trained features for all samples. Subsequently, in multiple iterations of active learning, a simple MLP classifier is trained using the pre-computed features for sample selection. After the sample selection, the entire model is fine-tuned once. However, we observe that, in comparison to the standard active learning method (fine-tuning the whole pre-trained model in each active learning iteration), SVPp may compromise the effectiveness of active learning, resulting in additional annotation costs. As depicted in fig. 1, we noticed that while SVPp significantly reduces active learning time (training costs), its decline in active learning performance leads to a notable increase in annotation costs. Practitioners therefore have to face the dilemma of weighing computational efficiency against overall cost (the sum of labeling cost and the training cost).
@@
In light of this challenge, in this paper, we analyze the factors contributing to the decline in active learning performance when using SVPp in sec. 4. Beyond the intuition-aligned reason for SVPp active learning performance decline: fine-tuned features being more capable of distinguishing sample categories than pre-trained features, leading to the selection of redundant samples (i.e., samples the fine-tuned model can already predict correctly), our analysis reveals an intriguing discovery: when pre-trained features are of comparable quality to fine-tuned features, not all sample selection differences result in a decline in SVPp performance. Moreover, some performance drops are due to samples selected by SVPp causing substantial modifications to the pre-trained model during fine-tuning.
@@
Based on these insights, we propose an aligned selection via proxy strategy, ASVP, to enhance the performance of efficient active learning in sec. 5. First, we align the pre-computed features used by the proxy model with the one used in standard active learning. Specifically, we update the pre-computed features based on the fine-tuned model when the pre-trained features’ ability to distinguish sample categories significantly lags behind the fine-tuned features. In addition, we switch the training method between linear-probing then fine-tuning (LP-FT) (Kumar et al., 2022) and fine-tuning (FT) to mitigate the impact of samples selected by the proxy model on the model’s performance.
@@
Extensive experimental validation shows that our method improves the effectiveness of efficient active learning, achieving similar or better total cost savings compared to those of standard active learning, while still maintaining computational efficiency.
@@
The contributions of this paper can be summarized as follows: (1) This paper empirically analyzes which sample selection differences contribute to the performance drop when employing SVPp. (2) We propose a novel efficient active learning approach, aligned selection via proxy (ASVP), based on the detailed analysis. It improves the performance of efficient active learning while incurring a marginal amount of computation time. (3) We introduce a novel and practical evaluation metric for efficient active learning: the sample saving ratio. This metric directly quantifies the savings in annotation achieved by employing active learning strategies compared with the random baseline. It also facilitates the estimation of overall savings, including savings in both computational and labeling costs, thereby assisting practitioners in making a decision on whether to use efficient active learning strategies.
@@ 2 Related Work
Active learning is a technique that minimizes the number of labels needed for model training by selectively requesting annotations for samples. Numerous active learning strategies have been proposed, primarily including uncertainty sampling (Lewis & Catlett, 1994; Scheffer et al., 2001; Gal et al., 2017; Kirsch et al., 2019; Woo, 2022), diversity sampling (feature space coverage) (Sener & Savarese, 2018; Mahmood et al., 2022), their combinations (Yang et al., 2017; Ash et al., 2020; Shui et al., 2020), and some learning-based approaches (Sinha et al., 2019; Yoo & Kweon, 2019; Tran et al., 2019). With the advancement of large-scale unsupervised pre-training and fine-tuning, an increasing number of researchers have turned their attention to active fine-tuning, which fine-tuning the pre-trained model with samples selected by active learning strategies. Given the success of the pre-training fine-tuning mode, most works of active fine-tuning have focused on developing active learning strategies that can work with extremely limited annotations (Hacohen et al., 2022; Yehuda et al., 2022; Wen et al., 2023). Furthermore, researchers have validated the performance of existing active learning methods when applied to pre-training fine-tuning mode, demonstrating that traditional margin sampling exhibits strong performance (Zhang et al., 2024; Emam et al., 2021).
Efficient active learning. Most active learning methods operate in a multi-round mode, iteratively training models and selecting samples to label. This mode needs significant training time and costs. In the context of active fine-tuning, this phenomenon becomes more pronounced, as larger models are better equipped to effectively leverage data in unsupervised pre-training (Chen et al., 2020b; Oquab et al., 2023).
Increasing the active learning batch size, which refers to the number of samples selected per active learning iteration, is one solution to enhance computational efficiency (Citovsky et al., 2021; Zhang et al., 2022). However, it weakens the performance of active learning, and there remains the considerable computational cost associated with training the entire network multiple times. Therefore, recent research has explored single-shot active learning based on pre-trained features (Xie et al., 2023). However, as the number of labels increases, the effectiveness of this method gradually diminishes in comparison to other standard active learning strategies.
Another solution to boost active learning efficiency is to use less computationally intensive models or training methods as proxy tasks for sample selection, known as Selection via Proxy (SVP) (Coleman et al., 2019). Typical proxy tasks include reducing model complexity, such as using models like ResNet-8 or ResNet-14, early stopping, or training linear classifiers or MLPs based on pre-trained features (Zhang et al., 2024). However, these proxy tasks often have an impact on active learning performance and may result in savings in computation costs not outweighing the increase in annotation costs. As a result, practitioners are often faced with a trade-off between computational efficiency and the overall cost.
@@ 3 Preliminary
3.1 Selection via Proxy based on Pre-computed Features
We briefly review the Selection via Proxy framework based on pre-computed features, SVPp, as shown in fig. 3. We denote the labeled set as $L$, the unlabeled set as $U$. The pre-trained neural network backbone is defined as $f(\cdot;\theta_{p}):\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $\theta_{p}$ represents weights of the pre-trained model, $\mathcal{X}$ is the data space and $\mathbb{R}^{d}$ is feature space. And we define the predictor as $h(\cdot;\theta_{h}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, where $\theta_{h}$ is weights of the predictor and $\mathbb{R}^{c}$ is the output space. SVPp follows a three-step process. First, all samples are forward-passed through $f(\cdot;\theta_{p})$, and the pre-trained features are saved. Subsequently, the proxy model is trained, and samples are selected iteratively based on the active learning strategy. In this process, the predictor $h(\cdot;\theta_{h})$ serves as the proxy model, utilizing pre-computed features as its input. The final step is fine-tuning the pre-trained model $h(f(\cdot;\theta_{p});\theta_{h})$ based on the obtained labels.
@@ 3.1 Selection via Proxy based on Pre-computed Features
3.2 LogME-PED
LogME-PED is an efficient metric for selecting pre-trained models, which outputs a score reflecting the performance of fine-tuned models. LogME uses evidence as a metric to assess the compatibility between a given pre-trained feature and the labels of the training samples (You et al., 2021), as shown in equation 1, where $F$ denotes the pre-trained features of the labeled set and $w$ denotes the parameters of the linear classifier. For specific calculation methods, refer to (You et al., 2021).
[…]
PED is a physically inspired model, that simulates feature learning dynamics through multiple iterations (Li et al., 2023). The approach draws an analogy between the optimization objective during fine-tuning and the concept of potential energy in physics. In fine-tuning, adjusting the model parameters forces the features of different classes to be distinguished, akin to the movement of objects influenced by forces in physics. Following this intuition, PED models the features of different classes as separate small balls, where the mean and standard deviation of each class’s features correspond to the position and radius of the ball, respectively. The force between balls is proportional to their overlapping length. In this physical model, the positions and forces of the objects are updated iteratively until convergence. Finally, the updated features are regarded as the features after fine-tuning the pre-trained model. Then, LogME is computed based on these updated features to assess the performance of the fine-tuned model.
@@ 4 Observation and Analysis
Since practitioners utilizing proxy models for efficient active learning aim to reduce sample selection time while minimizing the loss of active learning gains (between standard active learning based on fine-tuned models and SVPp), the exact overlap between samples selected by the proxy model and those selected by the fine-tuned model is not their primary concern. Therefore, in this section, we first investigate which differences in sample selection lead to the degradation of SVPp active learning performance.
4.1 Impact of Sample Selection Differences on SVPp Performance
We categorize the samples selected by both the proxy model and the fine-tuned model into different regions based on whether the models can correctly predict the samples and the active learning metrics (such as margin (Scheffer et al., 2001), entropy (Lewis & Catlett, 1994), etc.). As shown in fig. 4, the x and y axes represent whether the proxy model and the fine-tuned model can correctly predict the samples, respectively. The closer a sample is to the origin, the more likely it is to be selected in active learning. For uncertainty-based active learning strategies (e.g., margin, entropy, confidence (Wang & Shang, 2014)) and strategies combining uncertainty and diversity (e.g., BADGE (Ash et al., 2020)), a smaller distance from the origin indicates higher uncertainty, such as a smaller margin predicted by the model. For distance-based active learning strategies (e.g., coreset (Sener & Savarese, 2018)), a smaller distance from the origin indicates a greater distance between the sample and the labeled set. Most effective active learning strategies for fine-tuning pre-trained models are uncertainty-based or combine uncertainty with diversity (where the latter extends the former by selecting diverse samples from the uncertainty candidates) (Zhang et al., 2024; Emam et al., 2021), this paper will use uncertainty-based active learning as an example for analysis.
Standard active learning based on the fine-tuned model selects samples from regions O, A, and B, while efficient active learning based on the proxy model selects samples from regions O, C, and D. In other words, compared to the fine-tuned model, the proxy model misses samples from regions A1, A2, B1, and B2 while additionally selecting samples from regions C and D.
@@ 4.1 Impact of Sample Selection Differences on SVPp Performance
Figure 4: Sample Selection Discrepancy: Proxy vs. Fine-tuned Models. For instance, in uncertainty-based active learning strategies, the two axes represent the confidence of predictions for the proxy model and the fine-tuned model, with the positive half-axis indicating correct predictions and the negative half-axis indicating incorrect predictions. AL strategy selects samples with the lowest prediction confidence, meaning the proxy model chooses samples from regions O, C, and D, while the fine-tuned model selects samples from regions O, A, and B. Depending on whether the proxy model correctly predicts the samples, regions A and B are further divided into subregions A1, A2, B1, and B2. Both the proxy model and the fine-tuned model confidently predict the remaining white region, hence they do not select samples from it.
Empirical Evidence
We conducted replacement experiments to demonstrate the impact of missing samples from regions A1, A2, B1, and B2 on SVPp active learning performance. In this section, we used one of the state-of-the-art strategies, margin sampling, with fine-tuned pre-trained models. A total of 400k samples were selected over 16 rounds, with the first 10k samples selected randomly. The experiments were conducted on ImageNet, with detailed settings provided in sec. 6.1.
Figure 5: The impact of different regions on SVPp active learning performance. Replacing samples selected by the proxy model with those from regions A1, A2, B1, B2. The active learning gain refers to the difference in accuracy between active learning and the random baseline.
Specifically, during each round of sample selection, we replaced the tail samples selected by the proxy model (i.e., those with the largest margins) with samples from regions A1, A2, B1, and B2. As shown in fig. 5, we observed that not all differences in sample selection caused performance degradation in SVPp active learning. Replacing samples with those from regions A2 and B2 improved SVPp active learning performance, replacing samples from region A1 had no effect, while replacing samples from region B1 decreased performance. Based on these observations, to improve the performance of SVPp active learning, we should focus on the sample selection differences from regions A2 and B2.
To investigate why samples from regions A1 and B1 fail to improve SVPp performance, we analyzed the difficulty of samples across different regions. Recent studies suggest that hard-to-learn samples are likely to undermine active learning performance (Karamcheti et al., 2021). A1 and B1 regions, compared to A2 and B2, likely contain a higher proportion of difficult samples. This is because the fine-tuned model’s predictions on these samples show high uncertainty, while the proxy model’s predictions on them are wrong. To validate this hypothesis, we utilized Dataset Maps (Swayamdipta et al., 2020) to assess sample difficulty. Dataset Maps depict sample learnability by analyzing training dynamics from models trained on the entire dataset, using metrics such as average confidence for the correct class and prediction accuracy across training epochs to quantify difficulty. Following  Karamcheti et al. (2021), we categorized sample difficulty based on mean confidence levels: very hard (<0.25), hard (>=0.25 and <0.5), medium (>=0.5 and <0.75), and easy (>=0.75). As illustrated in fig. 6, regions A1 and B1 indeed contain a higher proportion of very hard samples.
(a) A1
(b) A2
(c) B1
(d) B2
Figure 6: Difficulty distribution of samples from different regions.
4.2 Role of Region A2 and B2 in Improving SVPp Performance
Those missed samples (Region A2, B2) that can improve SVPp performance share a common characteristic: the proxy model can confidently predict them correctly, while the fine-tuned model shows high uncertainty in its predictions. This indicates that these samples are well-distinguished using pre-trained features, whereas they are harder to differentiate in the fine-tuned features. Including these samples in the next round of active learning will help the fine-tuned model distinguish them. A question arises: are the pre-trained feature manifold and the fine-tuned feature manifold (training with samples from regions A2 and B2) similar?
Empirical Evidence
To answer this question, we conducted a backbone exchange experiment to compare the similarity between the fine-tuned and pre-trained features. We put the linear classifier of the fine-tuned model on top of the pre-trained backbone and evaluated its accuracy loss on the test set. The greater the accuracy drop after exchanging the backbone, the more dissimilar the fine-tuned features are from the pre-trained features.
We created five sets of training samples by replacing 0% (w/o replacing), 5%, 10%, 15%, and 20% of the samples selected by the proxy model with the samples from regions A2 and B2 that had the smallest margins. After training on these sets, the accuracy loss is evaluated by putting the fine-tuned model’s linear classifier on top of the pre-trained backbone. As shown in fig. 8, the accuracy drop decreased as more samples from A2 and B2 were included in the train set, indicating that incorporating these samples helps maintain the feature manifold similar to the pre-trained model.
Figure 7: The influence of samples from regions A2 and B2 on the similarity between fine-tuned and pre-trained features. Exchange-backbone accuracy loss refers to the accuracy difference between the fine-tuned model’s linear classifier placed on the pre-trained backbone and the fine-tuned model.
Figure 8: Comparison of active learning performance based on the fine-tuned model (standard active learning) and proxy model when employing different training methods. The active learning gain refers to the difference in accuracy between active learning and the random baseline (using fine-tuning).
Based on this observation, another intriguing question arises: can directly maintaining the similarity between fine-tuned and pre-trained features mitigate the performance drop caused by the proxy model missing samples from regions A2 and B2?
Empirical Evidence
To explore this, we employed two training methods: fine-tuning and LP-FT (Linear Probing and then Fine-Tuning), to compare the performance differences between SVPp (sample selection via proxy model and evaluation via fine-tuning or LP-FT) and standard active learning (both sample selection and evaluation via fine-tuning or LP-FT). In LP-FT, linear probing is followed by fine-tuning to preserve the similarity between the fine-tuned features and the pre-trained features.
As shown in fig. 8, after applying LP-FT, the active learning performance based on the proxy model was almost identical to that based on the LP-FT model. This suggests that samples from regions A2 and B2 enhance active learning performance by preserving the similarity between the fine-tuned and pre-trained feature manifold. Therefore, directly preserving this similarity compensates for the decrease in active learning performance caused by the proxy model’s omission of samples from regions A2 and B2.
4.3 Discussions
The previous section discussed that not all differences in sample selection between SVPp and standard active learning result in performance differences (only those from regions A2 and B2 do). Moreover, when using LP-FT to train the model, the sample selection differences from regions A2 and B2 do not cause a decline in SVPp active learning performance. Next, we discuss the limitations of the above analysis.
Firstly, given the training samples, we always aim to choose the training method that achieves the highest accuracy. Therefore, it is only meaningful to use LP-FT to compensate for the proxy model missing samples from regions A2 and B2 if LP-FT can achieve accuracy that is at least as good as fine-tuning. Secondly, as the number of selected samples increases, the performance gap between the fine-tuned model and the proxy model may widen (Cai et al., 2021). This could reduce the active learning gain of SVPp, thus affecting the previous conclusion about the impact of missed regions (A1, A2, B2, B2) by the proxy model on active learning performance. Therefore, we further analyzed the proportion of samples from regions C and D, as well as their impact on active learning gain.
(a) Proportion of Region C and D
(b) Active Learning Gain
Figure 9: Observations on region C and D: (a) the proportion of region C and D, (b) comparison of active learning gain when selecting only samples from region C or D. The active learning gain refers to the difference in accuracy between active learning and the random baseline.
As shown in fig. 9(a), with more samples selected during active learning, the proportion of samples from region C selected by the proxy model increases, while the proportion from region D decreases. This aligns with intuition: as the number of annotations increases, the features of the fine-tuned model become better at distinguishing between sample categories than those of the proxy model, increasing correctly predicted samples (samples from region C). Additionally, as shown in fig. 9(b), selecting samples from region C results in an active learning gain of less than 0, meaning it performs worse than random selection. These two observations imply that as more samples are selected during active learning, the proxy model selects more samples from region C, causing its active learning performance to decline. As the proxy model’s active learning performance declines, it may lead to all missed regions (A1, A2, B1, B2) contributing to performance differences between SVPp and standard active learning.
In summary, when LP-FT training achieves performance comparable to fine-tuning and the proportion of samples from region C is small, using LP-FT can prevent the decline in active learning gains caused by SVPp. The next section discusses how to adjust proxy model-based active learning to build an efficient active learning framework when these conditions are not met.